OpenStudy (anonymous):

3/5a^4(20a^6-10a^3-5/7) Someone help please

OpenStudy (mertsj):

\[\frac{3}{5}a^4(20a^6-10a^3-\frac{5}{7})\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Use the distributive property

OpenStudy (mertsj):

First multiply: \[\frac{3}{5}a^4\times 20a^6\]

OpenStudy (mertsj):

What do you get for that?

OpenStudy (anonymous):

12a^10

OpenStudy (mertsj):

Very good.

OpenStudy (mertsj):

Now do: \[\frac{3}{5}a^4 \times -10a^3\]

OpenStudy (anonymous):

-6a^7

OpenStudy (mertsj):

And finally: \[\frac{3}{5}a^6 \times - \frac{5}{7}\]

OpenStudy (anonymous):

-15/35a^4

OpenStudy (mertsj):

Cancel the 5's before you multiply

OpenStudy (anonymous):

3/7a^4

OpenStudy (mertsj):

Make it negative

OpenStudy (mertsj):

Now write your answer:

OpenStudy (anonymous):

\[12a ^{10}-6a ^{7}-\frac{ 3 }{ 7 }a ^{4}\]

OpenStudy (mertsj):

\[12a ^{10} - 6a^7-\frac{3}{7}a^4\]

OpenStudy (mertsj):

Perfect.

OpenStudy (anonymous):

can you help me with another

OpenStudy (mertsj):

Possibly

OpenStudy (anonymous):

(7a+1)(9a+3)

OpenStudy (anonymous):

its factoring

OpenStudy (anonymous):

I got 63a^2+30+3

OpenStudy (anonymous):

what did i do wrong?

OpenStudy (mertsj):

The middle term should be 21a + 9a which is 30a

OpenStudy (mertsj):

But you said it is factoring and you multiplied the two factors together so what do the directions say to do?

OpenStudy (anonymous):

I need to Foil

OpenStudy (anonymous):

i meant foiling not factoring

OpenStudy (mertsj):

ok. So it should be 63a^2 + 30a +3

OpenStudy (anonymous):

i was close enough

OpenStudy (mertsj):

Except what you had was really 63a^2+33

OpenStudy (anonymous):

(8x-7)(x+5) is the answer 8x^2+33x-35